# Effective Quantization of Muon Optimizer States

### v1

@@ 3.2 The Muon algorithm
where $m$ and $n$ are the dimensions of $\mathbf{M}$. Liu et al. (2025) claim that with Eq 2, hyperparameters such as learning rate and weight decay can be shared across matrix and non-matrix parameters. In the rest of the paper, any mention of Muon refers to the version in Equation 2, unless stated otherwise. It is important to note that any non-matrix parameters and input/output parameters are optimized using AdamW, leaving Muon to focus on matrix-valued hidden parameters. Complete algorithms can be found in Appendix B.
4 Methods
@@ 4.1 Why is 8-bit AdamW unstable under linear quantization?
Dettmers et al. (2021) observed that naïve 8-bit linear quantization performs poorly because it allocates too little resolution to small-magnitude entries, yielding large relative errors precisely where optimizer states concentrate most of their mass. We make this behavior concrete in Theorem 1 by analyzing how quantization error propagates and is magnified through the second-moment accumulator $\mathbf{v}$. In particular, we show that if moderate-size gradient coordinates occur with non-negligible probability, then the expected squared error of one step of Adam with linearly quantized states diverges as the numerical stabilizer $\epsilon\to 0$. The result holds for standard Adam hyperparameters; constants are kept explicit to emphasize practical regimes (e.g., $\epsilon\approx 10^{-8}$) where the error is already proved to be orders of magnitude larger than the unquantized update norm.
Let $Q$ denote the 8-bit linear quantization operator from Definition 1. We analyze the base Adam algorithm without weight decay [^3] The result extends immediately to AdamW, since the decay of $\bm{\theta}^{(0)}$ affects both $\tilde{\bm{\theta}}^{(1)}$ and $\bm{\theta}^{(1)}$ equally and therefore cancels out.. The quantized variant applies $Q$ to the moment estimates before forming the update. All algorithmic details, definitions, and proofs are provided in Appendix C.
@@
Let $\bm{\theta}^{(1)}$ denote the parameters after one step of Adam as given in Algorithm 4, and let $\tilde{\bm{\theta}}^{(1)}$ denote the parameters after one step of the same algorithm with 8-bit linear quantization applied to the moment estimates (Definition 1), i.e.:
$\tilde{\bm{\theta}}^{(1)}=\bm{\theta}^{(0)}-\alpha\cdot\frac{Q(\mathbf{m}^{(1)})}{\sqrt{Q(\mathbf{v}^{(1)})}+\epsilon}.$
Suppose that each entry of $\mathbf{g}^{(1)}\in\mathbb{R}^{d}$ satisfies $\mathbb{P}\!\left(\tfrac{\|\mathbf{g}\|_{\infty}}{60}<|\mathbf{g}_{i}|<\tfrac{\|\mathbf{g}\|_{\infty}}{16}\right)\geq\nu$ and $\|\mathbf{g}\|_{\infty}\geq g_{\infty}>256\,\epsilon$ with probability one. Then
$\mathbb{E}\,\|\bm{\theta}^{(1)}-\tilde{\bm{\theta}}^{(1)}\|_{2}^{2}\;\geq\;\frac{d\nu\,\alpha^{2}g_{\infty}^{2}}{(256\,\epsilon)^{2}}.$
@@ 4.2 The curious case of 8-bit SGD with momentum
$\tilde{\bm{\theta}}^{(t+1)}=\bm{\theta}^{(t)}-\eta(\mathbf{g}^{(t)}+\rho Q(\mathbf{m}^{(t)}))~~\text{and}~~\bm{\theta}^{(t+1)}=\bm{\theta}^{(t)}-\eta(\mathbf{g}^{(t)}+\rho\mathbf{m}^{(t)}).$
From any point $\bm{\theta}^{(t)}$ and any momentum state $\mathbf{m}^{(t)}$, if $Q$ is as in Definition 1, then
$\|\tilde{\bm{\theta}}^{(t+1)}-\bm{\theta}^{(t+1)}\|_{2}^{2}\leq d\eta^{2}\rho^{2}\left(\frac{\|\mathbf{m}^{(t)}\|_{\infty}}{127}\right)^{2}.$
Empirically, we confirm this difference. We train a ResNet-50 model (He et al., 2016) on the ImageNet dataset (Deng et al., 2009), using a standard training regime of 90 epochs. We compare AdamW and variants of SGD with momentum. Results are in Table 2. Surprisingly, SGD with linear 8-bit quantization achieves the same high validation top-1 accuracy (Goyal et al., 2017) of 76%+ as full-precision SGD. AdamW underperforms when compared to SGD (a well-known result on image classification training), while AdamW with linear quantization diverges immediately. See Appendix A.1.1 for extended details.
@@
Table 2: Top-1 validation accuracy (%) after 90 epochs for SGD+M and AdamW in FP32 and with 8-bit linear quantization. “—” indicates that Adam with linear quantization diverged.
Together, these theoretical and empirical results demonstrate that the instability of quantized Adam is specifically driven by the quantization of the second-moment term in the denominator.
4.3 The 8-bit Muon algorithm
Since SGD works well with linear quantization, we now propose 8-bit quantized variants for the Muon algorithm that leverage either linear or dynamic schemes. In Table 1, we introduce various variants of Muon and AdamW that leverage different quantization schemes. Since Muon uses orthogonalization only for hidden matrix-valued parameters, the other parameters like embeddings and classifier heads are usually optimized with AdamW. The variants listed in Table 1 consist of linear, dynamic, and even hybrid versions.
The 8-bit Muon update can be written as:
$\widetilde{\mathbf{M}}^{(t-1)}$ $=\mathrm{DQ}^{\text{mode}}_{B}\!\big(\mathbf{Z}^{(t-1)},\mathcal{S}^{(t-1)}\big),$ (3)
$\mathbf{M}^{(t)}$ $=\beta\,\widetilde{\mathbf{M}}^{(t-1)}+\nabla f_{t}\!\big(\mathbf{W}^{(t-1)}\big),$
$\mathbf{U}^{(t)}$ $=\mathrm{NS}\!\big(\mathbf{M}^{(t)}\big),$
$\mathbf{W}^{(t)}$ $=(1-\alpha\lambda)\,\mathbf{W}^{(t-1)}\;-02\,\alpha\,\sqrt{\max(m,n)}\,\mathbf{U}^{(t)},$
$\big(\mathbf{Z}^{(t)},\mathbf{\mathcal{S}}^{(t)}\big)$ $=\mathrm{Q}^{\text{mode}}_{B}\!\big(\mathbf{M}^{(t)}\big).$
where $\mathbf{Z}$ refers to the compressed momentum buffer and $\mathbf{\mathcal{S}}$ is the associated state required to dequantize it. $\mathrm{Q}^{\text{mode}}_{B}$ is the quantization function, and $\mathrm{DQ}^{\text{mode}}_{B}$ dequantizes the compressed momentum vector for use in the update. Rest of the notation is borrowed from Equation 2. The complete algorithms for 8-bit Muon can be found in Appendix B.
We now claim that like SGD, even Muon admits a uniform bound under linear quantization. In fact, in Theorem 3 we show the following: when Muon uses an exact orthogonalization procedure (via the SVD), the quantization error bound for a single layer matches the SGD case up to an additional dependence on the smallest singular value, $s$, of the momentum matrix. This dependence is natural, since the conditioning of the momentum controls the stability of the orthogonal factor. In practice, the margin $s$ is typically not small, so the guarantee has the same qualitative form as the SGD result.

### v3

@@ 3.2 The Muon Algorithm
where $m$ and $n$ are the dimensions of $\mathbf{M}$. Liu et al. (2025) claim that with eqn. (2), hyperparameters such as learning rate and weight decay can be shared across matrix and non-matrix parameters. In the rest of the paper, any mention of Muon refers to the version in eqn. (2), unless stated otherwise. It is important to note that any non-matrix parameters and input/output parameters are optimized using AdamW, leaving Muon to focus on matrix-valued hidden parameters.
4 Experiments
This section presents a thorough empirical study of using linear and/or dynamic quantization to improve the memory efficiency of the Muon optimizer with minimal quality degradation. Recall that Muon is applied only to the two-dimensional hidden-layer weight matrices, while AdamW is used for all remaining parameters; see Table 1 for details.
4.1 The Quantized Muon Algorithm
Throughout our experiments, we use the Muon variant of (Liu et al., 2025), while compressing the optimizer state stored between iterations via linear or dynamic blockwise 8-bit quantization.
$\widetilde{\mathbf{M}}^{(t-1)}$ $=\operatorname{Dequantize}_{\mathcal{S}}\big(\mathbf{Z}^{(t-1)}\big),$ (3)
$\mathbf{M}^{(t)}$ $=\beta\,\widetilde{\mathbf{M}}^{(t-1)}+\nabla f\!\big(\mathbf{W}^{(t-1)}\big),$
$\mathbf{O}^{(t)}$ $=\mathrm{NS}\!\big(\mathbf{M}^{(t)}\big),$
$\mathbf{W}^{(t)}$ $=(1-\alpha\lambda)\,\mathbf{W}^{(t-1)}\;-02\alpha\sqrt{\max(m,n)}\mathbf{O}^{(t)},$
$\mathbf{Z}^{(t)}$ $=\operatorname{Quantize}_{\mathcal{S}}\big(\mathbf{M}^{(t)}\big).$
where $\mathbf{Z}$ denotes the compressed momentum buffer and $\mathcal{S}$ denotes the associated auxiliary state used for dequantization. The remaining notation follows Equation 2.
4.2 Pre-training with 8-bit Muon
Architectures. For pre-training, we train a modified GPT-2 architecture (Radford et al., 2019) from scratch, replacing learned absolute positional embeddings with rotary positional embeddings (RoPE) (Su et al., 2024). To study scaling, we consider five model sizes: Small (162M), Medium (405M), Large (834M), XL (1.4B), and XXL (2.7B). For brevity, we refer to this architecture family as GPT throughout the paper. Additional architectural details are provided in Table 7 in Appendix A.1.2.
Datasets. We pre-train on FineWeb-Edu (Penedo et al., 2024). Following the compute-optimal scaling guidance of Hoffmann et al. (2022), we use approximately 20 tokens per parameter, yielding a training set size that increases with model size (the Chinchilla-optimal setting). For the 2.7B model, this corresponds to roughly 54B FineWeb tokens. We evaluate validation loss on 150K samples from the FineWeb validation split, totaling approximately 300M tokens.
Training details. We pre-train the GPT models from scratch on FineWeb. For AdamW, we use $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. For Muon, we set the momentum parameter to $0.95$. A complete list of hyperparameters and training configurations can be found in Appendix A.1.2. For quantized AdamW and Muon variants, we use a block size of 2048. We fix the decoupled weight decay to 0.1 in all experiments.
We adopt the warmup-stable-decay (WSD) learning rate schedule (Hu et al., 2024), with linear warmup from 0 to the peak learning rate over the first 10% of training steps, and linear decay to zero over the final 10% of steps. For each model size, we use the peak learning rate and global batch size recommended by Brown et al. (2020), which were tuned for GPT-style models trained using AdamW. For Muon, we do not tune the peak learning rate; instead, we reuse the AdamW peak learning rates, consistent with our goal that Muon should serve as a drop-in replacement for AdamW without requiring learning-rate or weight-decay re-tuning (Liu et al., 2025).
Evaluation criteria Our evaluation of pre-training is two-pronged (a) validation loss and (b) benchmark performance on six different tasks using the lm-eval harness (Gao et al., 2024). These tasks include MMLU (Hendrycks et al., 2020), LAMBADA (Paperno et al., 2016), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), ARC-Challenge and ARC-Easy (Clark et al., 2018).
4.2.1 Results
Figure 1: Validation loss for GPT-2 Medium across 7 variants.
Medium-scale ablation: stability and sensitivity to quantization.
We first perform a controlled ablation on GPT-Medium, sweeping all stable 8-bit Muon variants from Table 1 and comparing them against AdamW baselines. Figure 1 plots validation loss versus training step. Under this setup, AdamW-8L diverges early (confirmed in separate runs and therefore omitted from the figure), consistent with the observations of Dettmers et al. (2021).
The AdamW baselines lag behind throughout training, with AdamW-8D consistently the worst among the plotted methods and AdamW-32 also trailing the Muon family. Secondly, all Muon-based variants—including hybrids that quantize the AdamW-associated states and variants that use either dynamic or linear quantization for Muon-associated states—track Muon-32 extremely closely, with near-overlapping trajectories across the full training horizon.
Scaling results: pretraining validation loss.
Having established stability and quantization robustness at the Medium scale, we next scale to a broader sweep over model sizes. We compare the full-precision Muon-32 baseline against two practically relevant 8-bit variants, Muon-8L/AdamW-32 and Muon-8L/AdamW-8D. We focus on these Muon-8L-based implementations because they require minimal engineering overhead, making them straightforward to integrate into distributed training frameworks (e.g., (Ahn et al., 2025)).
Table 2 reports the final pretraining validation loss for Small through XXL, aggregated over five seeds; the full training and validation curves are provided in Figures 3 and 4 in Appendix B.6. Across all sizes, Muon-8L/AdamW-32 matches full-precision Muon-32 up to run-to-run variability: the mean loss differs by at most $0.002$ (e.g., Small: $3.120$ vs. $3.122$; XXL: $2.345$ vs. $2.346$). Meanwhile, Table 3 of Dettmers et al. (2021) implies a loss gap of $0.006$ between 32-bit Adam and 8-bit Adam with block-wise dynamic quantization, both (i) with stable embeddings and (ii) without stable embeddings [^2] We convert perplexity to loss via $\text{loss}=\ln(\text{Perplexity})$. Muon-8L/AdamW-8D remains competitive but shows a small and consistent degradation relative to Muon-32, with the largest gap at XL ($2.495$ vs. $2.508$). Overall, these results demonstrate that quantizing the Muon-associated state to 8-bit while keeping the AdamW component in full precision (Muon-8L/AdamW-32) preserves the optimization behavior of Muon-32 across five model scales, while delivering the memory benefits of 8-bit state compression.
Downstream generalization
We next assess generalization by evaluating each pretrained model on six established benchmarks. As shown in Table 3, hybrids that keep the AdamW-associated states in full precision (e.g., Muon-8L/AdamW-32) remain on par with the full-precision Muon-32 baseline across model sizes, with differences in the six-task average that are within seed-level variation. In contrast, the fully 8-bit hybrid Muon-8L/AdamW-8D exhibits a small but consistent degradation in the averaged downstream score, most noticeably at Medium, XL and XXL, mirroring its slightly worse pretraining validation loss. These results suggest that aggressively quantizing the AdamW component can be the limiting factor for downstream transfer, whereas quantizing the Muon-associated state alone preserves generalization.
Takeaway.
Across a targeted ablation (Figure 1) and a scale sweep (Tables 23), we find that (i) Muon-based 8-bit variants train reliably; and (ii) Muon-8L matches full-precision Muon-32 in both pretraining validation loss and downstream generalization across model sizes. These findings support Muon-8L/AdamW-32 as a strong default for memory-efficient pretraining without sacrificing quality.
Model size Optimizer Validation loss
Small Muon-32 $3.120_{\scriptscriptstyle 0.002}$
Muon-8L/AdamW-32 $3.122_{\scriptscriptstyle 0.002}$
Muon-8L/AdamW-8D $3.129_{\scriptscriptstyle 0.001}$
Medium Muon-32 $2.819_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-32 $2.820_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-8D $2.824_{\scriptscriptstyle 0.001}$
Large Muon-32 $2.620_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-32 $2.621_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-8D $2.623_{\scriptscriptstyle 0.001}$
XL Muon-32 $2.495_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-32 $2.496_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-8D $2.508_{\scriptscriptstyle 0.001}$
XXL Muon-32 $2.345_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-32 $2.346_{\scriptscriptstyle 0.001}$
Muon-8L/AdamW-8D $2.352_{\scriptscriptstyle 0.001}$
Table 2: Pretraining validation loss across 5 seeds, reported as mean with standard deviation.
Model size Optimizer ARC-C ARC-E BoolQ HellaSwag LAMBADA MMLU Avg. (6 tasks)
Small Muon-32 $0.204_{\scriptscriptstyle 0.008}$ $0.503_{\scriptscriptstyle 0.004}$ $0.571_{\scriptscriptstyle 0.028}$ $0.285_{\scriptscriptstyle 0.002}$ $0.203_{\scriptscriptstyle 0.006}$ $0.230_{\scriptscriptstyle 0.001}$ 0.332
Muon-8L/AdamW-32 $0.211_{\scriptscriptstyle 0.003}$ $0.505_{\scriptscriptstyle 0.007}$ $0.574_{\scriptscriptstyle 0.031}$ $85_{\scriptscriptstyle 0.001} 0.202$$02_{\scriptscriptstyle 0.008}$ $0.230_{\scriptscriptstyle 0.001}$ 0.335
Muon-8L/AdamW-8D $0.205_{\scriptscriptstyle 0.009}$ $0.506_{\scriptscriptstyle 0.008}$ $0.580_{\scriptscriptstyle 0.023}$ $83_{\scriptscriptstyle 0.001} 0.202$$02_{\scriptscriptstyle 0.006}$ $0.230_{\scriptscriptstyle 0.000}$ 0.334
Medium Muon-32 $0.236_{\scriptscriptstyle 0.005}$ $0.564_{\scriptscriptstyle 0.009}$ $0.606_{\scriptscriptstyle 0.010}$ $0.316_{\scriptscriptstyle 0.001}$ $0.272_{\scriptscriptstyle 0.004}$ $0.231_{\scriptscriptstyle 0.004}$ 0.371
Muon-8L/AdamW-32 $0.247_{\scriptscriptstyle 0.009}$ $0.563_{\scriptscriptstyle 0.004}$ $0.613_{\scriptscriptstyle 0.007}$ $0.315_{\scriptscriptstyle 0.002}$ $0.274_{\scriptscriptstyle 0.006}$ $0.232_{\scriptscriptstyle 0.003}$ 0.374
Muon-8L/AdamW-8D $0.235_{\scriptscriptstyle 0.008}$ $0.559_{\scriptscriptstyle 0.010}$ $0.595_{\scriptscriptstyle 0.011}$ $0.314_{\scriptscriptstyle 0.002}$ $0.272_{\scriptscriptstyle 0.006}$ $0.230_{\scriptscriptstyle 0.002}$ 0.368
Large Muon-32 $0.281_{\scriptscriptstyle 0.008}$ $20_{\scriptscriptstyle 0.014} 0.606$$06_{\scriptscriptstyle 0.012}$ $0.355_{\scriptscriptstyle 0.002}$ $0.342_{\scriptscriptstyle 0.005}$ $0.250_{\scriptscriptstyle 0.006}$ 0.409
Muon-8L/AdamW-32 $0.278_{\scriptscriptstyle 0.006}$ $0.626_{\scriptscriptstyle 0.007}$ $0.610_{\scriptscriptstyle 0.010}$ $0.356_{\scriptscriptstyle 0.002}$ $0.338_{\scriptscriptstyle 0.004}$ $0.251_{\scriptscriptstyle 0.009}$ 0.410
Muon-8L/AdamW-8D $0.272_{\scriptscriptstyle 0.008}$ $0.628_{\scriptscriptstyle 0.005}$ $0.611_{\scriptscriptstyle 0.007}$ $0.355_{\scriptscriptstyle 0.002}$ $0.335_{\scriptscriptstyle 0.007}$ $0.245_{\scriptscriptstyle 0.013}$ 0.408
XL Muon-32 $0.296_{\scriptscriptstyle 0.006}$ $0.653_{\scriptscriptstyle 0.012}$ $0.607_{\scriptscriptstyle 0.015}$ $0.382_{\scriptscriptstyle 0.002}$ $0.376_{\scriptscriptstyle 0.007}$ $0.255_{\scriptscriptstyle 0.006}$ 0.428
Muon-8L/AdamW-32 $0.298_{\scriptscriptstyle 0.004}$ $0.651_{\scriptscriptstyle 0.011}$ $0.603_{\scriptscriptstyle 0.017}$ $0.381_{\scriptscriptstyle 0.002}$ $0.373_{\scriptscriptstyle 0.003}$ $0.253_{\scriptscriptstyle 0.009}$ 0.427
Muon-8L/AdamW-8D $0.293_{\scriptscriptstyle 0.005}$ $0.649_{\scriptscriptstyle 0.007}$ $0.586_{\scriptscriptstyle 0.037}$ $0.379_{\scriptscriptstyle 0.001}$ $0.368_{\scriptscriptstyle 0.004}$ $0.254_{\scriptscriptstyle 0.006}$ 0.422
XXL Muon-32 $0.354_{\scriptscriptstyle 0.010}$ $0.713_{\scriptscriptstyle 0.003}$ $0.618_{\scriptscriptstyle 0.005}$ $0.428_{\scriptscriptstyle 0.001}$ $0.424_{\scriptscriptstyle 0.003}$ $0.261_{\scriptscriptstyle 0.006}$ 0.466
Muon-8L/AdamW-32 $0.352_{\scriptscriptstyle 0.004}$ $0.707_{\scriptscriptstyle 0.004}$ $0.618_{\scriptscriptstyle 0.008}$ $0.428_{\scriptscriptstyle 0.002}$ $0.422_{\scriptscriptstyle 0.008}$ $0.256_{\scriptscriptstyle 0.006}$ 0.464
Muon-8L/AdamW-8D $0.347_{\scriptscriptstyle 0.008}$ $0.706_{\scriptscriptstyle 0.013}$ $0.625_{\scriptscriptstyle 0.009}$ $0.425_{\scriptscriptstyle 0.002}$ $0.418_{\scriptscriptstyle 0.005}$ $0.260_{\scriptscriptstyle 0.010}$ 0.463
Table 3: Downstream performance across 5 seeds, reported as mean with standard deviation.
Comparison Win Rate (%) LC Win Rate (%)
Muon-32 vs AdamW-32 59.67 59.17
M-8L/A-32 vs AdamW-32 60.48 59.93
M-8L/A-32 vs Muon-32 50.75 50.37
M-8L/A-8D vs AdamW-32 58.47 57.91
M-8L/A-8D vs Muon-32 48.17 48.14
Table 4: Head-to-head AlpacaEval results on N=805N{=}805 prompts. Win Rate is the raw preference rate; LC Win Rate is length-controlled. Values are percentages; $50\%$ indicates a tie.
4.3 Fine-tuning with 8-bit Muon
Task Setup and Architecture We test the fine-tuning efficacy of the quantized Muon optimizer by using different variants to instruction-finetune the pre-trained models trained previously. Concretely, we choose two checkpoints of the GPT-XL model - one pre-trained with AdamW and the other pre-trained with Muon. It is empirically well-known that if a model is pre-trained with one optimizer, then empirically it is best to fine-tune it with the same optimizer (Liu et al., 2025; Team et al., 2025). Thus, we then proceed to finetune the AdamW-pretrained checkpoint with AdamW-32, and the Muon-pretrained checkpoint with Muon-32, Muon-8L/AdamW-32 , and Muon-8L/AdamW-8D.
Datasets We use the Alpaca instruction following dataset (Taori et al., 2023). The dataset contains about 52k samples for instruction following, consisting of a diverse array of topics from code generation to creative writing.
Training Details We use SFT on the GPT-XL model using the learning rate of $2\times 10^{-5}$. For each optimizer, For AdamW, we use $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. We fixed the weight decay to be $0.01$.
Evaluation Criteria We use AlpacaEval 2.0 (Dubois et al., 2024) (referred to hereinafter as AE2) as our main eval benchmark. AE2 is an automated evaluation framework that leverages a strong LLM-as-a-judge to compare the instruction following abilities of two different candidate LLMs on a variety of 805 prompts. AE2 reports both win rate and length-controlled (LC) win rate to mitigate verbosity bias. In our setup, we use GPT4 Turbo as the judge.
Once trained, we pit the four models against each other in a pairwise fashion to collect both win rate and LC win rate.
4.3.1 Results
Head-to-head AE2 results can be found in Table 4. The model fine-tuned with Muon-32 consistently outperforms the model fine-tuned with AdamW-32, achieving 59.67% win rate (59.17% LC win rate). Muon-8L/AdamW-32 performs similarly, achieving a large win over AdamW-32 and being tied in performance with Muon-32. Muon-8L/AdamW-8D also outperforms AdamW-32 and slightly underperforms when compared to Muon-32.
These findings mirror the results from pre-training, where we demonstrate that quantizing the Muon-associated part of the optimizer results in essentially no loss in model quality.
4.4 Memory footprint
Table 5 compares the persistent High Bandwidth Memory footprint of the optimizer state for variants, when profiled for the GPT XL and XXL models. For the XXL model, Muon-8D/AdamW-32 provides substantial relative savings of $\sim$79% and $\sim$62% when compared to AdamW-32 and Muon-32. Muon-8D/AdamW-8D is more aggressive, stretching the advantage to $\sim$86% and $\sim$75%.
For the smaller XL model, the savings are slightly smaller. This is because the size of the embedding and lm-head matrices remains the same across model sizes because of a fixed vocabulary.
[…]
Table 5: Optimizer states memory (GiB) for two model configurations (lower is better). M-8L(D)/A-8D yields the smallest optimizer-state footprint in both cases.
5 Understanding the Stability of Linear Quantization
Our experimental results in Section 4.1 emphasize the effectiveness of applying linear quantization to the Muon optimizer state. This is in stark contrast to AdamW, where linear quantization results in divergence and even sophisticated dynamic quantization schemes may show performance degradation. In this section, we determine the mechanism driving these observations.
@@ 5.1 Why is 8-bit AdamW Unstable under Linear Quantization?
Dettmers et al. (2021) observed that AdamW with naïve 8-bit linear quantization performs poorly because it allocates too little resolution to small-magnitude entries, yielding large relative errors precisely where optimizer states concentrate most of their mass. We make this behavior more precise in Theorem 1 by analyzing how quantization error propagates and is magnified through the second-moment accumulator $\mathbf{v}$. In particular, we show that if moderate-size gradient coordinates occur with non-negligible probability, then the expected squared error of one step of Adam with linearly quantized states diverges as the numerical stabilizer $\epsilon\to 0$. The result holds for standard Adam hyperparameters; constants are kept explicit to emphasize practical regimes (e.g., $\epsilon\approx 10^{-8}$) where the error is already proved to be orders of magnitude larger than the unquantized update norm.
Let $Q$ denote the 8-bit linear quantization operator (Definition 1). Note that our error anlaysis is formalized over real numbers, and so $Q$ composes the quantization and dequantization steps that perturb the optimizer state. We consider the base Adam algorithm without weight decay, though it extends to AdamW without loss of generality [^3] The result extends immediately to AdamW, since the decay of $\mathbf{w}^{(0)}$ affects both $\tilde{\mathbf{w}}^{(1)}$ and $\mathbf{w}^{(1)}$ equally and therefore cancels out.. All algorithmic details, definitions, and proofs are provided in Appendix B. Note that we will use lower case letters when referring to one-dimensional vectorized parameters and upper case when operating on the two-dimensional hidden state parameters addressed by Muon, e.g., $\mathbf{w}$ versus $\mathbf{W}$.
@@
Let $\mathbf{w}^{(1)}$ denote the parameters after one step of Adam as given in Algorithm 1, and let $\tilde{\mathbf{w}}^{(1)}$ denote the parameters after one step of the same algorithm with 8-bit linear quantization applied to the moment estimates (Definition 1), i.e.:
$\tilde{\mathbf{w}}^{(1)}=\mathbf{w}^{(0)}-\alpha\cdot\frac{Q(\mathbf{m}^{(1)})}{\sqrt{Q(\mathbf{v}^{(1)})}+\epsilon}.$
Suppose that each entry of gradient $\mathbf{g}^{(1)}\in\mathbb{R}^{d}$ satisfies $\mathbb{P}\!\left(\tfrac{\|\mathbf{g}\|_{\infty}}{60}<|\mathbf{g}_{i}|<\tfrac{\|\mathbf{g}\|_{\infty}}{16}\right)\geq\nu$ and $\|\mathbf{g}\|_{\infty}\geq g_{\infty}>256\,\epsilon$ with probability one. Then
$\mathbb{E}\,\|\mathbf{w}^{(1)}-\tilde{\mathbf{w}}^{(1)}\|_{2}^{2}\;\geq\;\frac{d\nu\,\alpha^{2}g_{\infty}^{2}}{(256\,\epsilon)^{2}}.$
We note that empirical validation of Theorem 1, as well as Theorems 2 and 3 below, can be found in Appendix B.5.
@@ 5.2 The Curious Case of 8-bit SGD with Momentum
$\tilde{\mathbf{w}}^{(t+1)}=\mathbf{w}^{(t)}-\eta(\mathbf{g}^{(t)}+\rho Q(\mathbf{m}^{(t)}))$
$~~\text{and}~~\mathbf{w}^{(t+1)}=\mathbf{w}^{(t)}-\eta(\mathbf{g}^{(t)}+\rho\mathbf{m}^{(t)}).$
From any point $\mathbf{w}^{(t)}$ and any momentum state $\mathbf{m}^{(t)}$, then with linear quantization, $Q$ (Definition 1),
$\|\tilde{\mathbf{w}}^{(t+1)}-\mathbf{w}^{(t+1)}\|_{2}^{2}\leq d\eta^{2}\rho^{2}\left(\frac{\|\mathbf{m}^{(t)}\|_{\infty}}{127}\right)^{2}.$
Empirically, we confirm this difference. We train a ResNet-50 model (He et al., 2016) on the ImageNet dataset (Deng et al., 2009), using a standard training regime of 90 epochs. We compare AdamW and variants of SGD with momentum. Results are in Table 6. We observe that SGD with linear 8-bit quantization achieves the same high validation top-1 accuracy (Goyal et al., 2017) of 76%+ as full-precision SGD. AdamW underperforms when compared to SGD (a well-known result on image classification training), while AdamW with linear quantization diverges immediately. See Appendix A.1.1 for extended details.
@@
Table 6: Top-1 validation accuracy (%) after 90 epochs for SGD+M and AdamW in FP32 and with 8-bit linear quantization. “—” indicates that AdamW with linear quantization diverged.
Together, these results emphasize that quantized Adam’s instability is driven specifically by quantizing the second-moment term in the denominator for moderately small coordinates.
5.3 Linear Quantization of the Muon Optimizer
Like SGD with momentum, the Muon optimizer only maintains a first order momentum state between iterations, avoiding the second moment normalization term in AdamW. Theorem 3 below shows that the Muon optimizer admits an analogous uniform error bound, theoretically explaining the stability of quantized Muon.
In Theorem 3 we show the following: when Muon uses an exact orthogonalization procedure (via the SVD), the quantization error bound for a single layer matches the SGD case up to an additional dependence on the smallest singular value, $s$, of the momentum matrix. This dependence is natural, since the conditioning of the momentum controls the stability of the orthogonal factor. Our analysis considers the exact polar factor to avoid introducing additional finite-iteration Newton–Schulz approximation error into the bound.